Week of 160208

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Alessandro (ATLAS), Charles (databases), Fernando (batch + grid), Julia (WLCG), Maarten (SCOD + ALICE), Maria A (WLCG)
  • remote: Aresh (PIC), Dario (ATLAS), David (FNAL), Dmytro (NDGF), Francesco (CNAF), Kyle (OSG), Michael (BNL), Oliver (CMS), Onno (NLT1), Pepe (PIC), Rolf (IN2P3), Tiju (RAL), Zoltan (LHCb)

Experiments round table:

  • ATLAS reports (raw view) -
    • Production continuing with MC and data reprocessing (some 2012 and 2015).
    • Deletion failures at CERN Saturday were due to incorrect settings of access protocol (changed on Friday), fixed on Sunday.
    • CREAM-CE problem related to Condor and Java update versions still not understood (coming and going). Still affecting several sites. No recent updates on GGUS:119258 .
      • Maarten: that ticket concerns a Condor bug in its handling of CREAM job leases
        • that issue is independent of the one due to the latest updates for Java 6, 7 and 8:
          • by default the use of MD5 gets disabled
          • adjust /usr/lib/jvm/jre/lib/security/java.security to re-enable it (and restart the affected services)
          • MD5 is used by the old CREAM client in condor < 8.3.x (it also does not support RFC proxies)
          • the SAM tests for ATLAS and CMS currently use condor-8.2.10 and can thus be affected
          • an upgrade to condor-8.4.4 is being looked into
          • that would still be for the old SAM-Nagios system on SL5
          • in a number of weeks the new ETF-Nagios system using condor-8.4.4 on SL6 should take over

  • CMS reports (raw view) -
    • CERN
      • Problem with creating VMs in Openstack solved, Stuck gridmanager daemon on the glideinWMS factory was the cause. Thanks for the help. GGUS:119312
    • to all sites:
      • CMS xrootd SAM test became mandatory (error instead of warning in case of failure) for all sites today, Monday, Feb 8th
    • Due to conflicts, I cannot participate in the meeting today. Sorry.

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Data Processing
      • Stripping 24 is ongoing
      • MC and User jobs
    • Site Issues
      • T0:
        • CERN EOS issue (GGUS:119410): Upload files to CERN-BUFFER
      • T1:

Sites / Services round table:

  • ASGC:
  • BNL: ntr
  • CNAF:
    • CVMFS issues on some WN are being investigated (LHCb ticket)
  • FNAL: ntr
  • GridPP:
  • IN2P3: ntr
  • JINR:
  • KISTI:
  • KIT:
  • NDGF: ntr
  • NL-T1:
    • A small reminder: in October SARA will move to a new datacenter. The experiments are advised to check whether they have single copy data at SARA and make sure they have replicas elsewhere when necessary.
      • Alessandro: how long will the downtime be?
      • Onno: it is foreseen to be 2 weeks
      • Alessandro: should we be concerned about potential data losses?
      • Onno: for such big moves there is a typical risk that 2% of the equipment may suffer HW problems afterwards
      • Maarten: that would be for computers and disks, while tape should - by design - be even less affected
      • Onno: will check if there exist estimates for tape problems
  • NRC-KI:
  • OSG: ntr
  • PIC:
    • outage the whole day of March 22 for the yearly electrical maintenance
  • RAL: ntr
  • TRIUMF:

  • CERN batch and grid services: ntr
  • CERN storage services:
  • Databases:
    • the firewall changes for the CMS databases will be done Wed afternoon
    • ALICE online DB maintenance went OK on Thu
  • GGUS:
  • Grid Monitoring:
  • MW Officer:

AOB:

Thursday

Attendance:

  • local: Charles (databases), Fernando (batch + grid), Luca (storage), Maarten (SCOD + ALICE)
  • remote: Andrew (NLT1), Chris (OSG), David C (ATLAS), David M (FNAL), Dmytro (NDGF), Kyle (OSG), Michael (BNL), Rolf (IN2P3), Sang-Un (KISTI), Tiju (RAL), Zoltan (LHCb)

Experiments round table:

  • ATLAS reports (raw view) -
    • (Re)-reprocessing of 2015 data started last night
    • Any news on CREAM-condor problem?
      • Monday's ticket was just closed by ATLAS as the matter seems under control
      • the SAM tests are still at risk, but so far there has not been any outcry

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Data Processing
      • Stripping 24 is almost at 100%, the merging productions are following slowly.
      • MC and User jobs
    • Site Issues
      • T0:
        • network router update happened yesterday (10 of Feb). we have not received any notification about the start and finish of the intervention.
          • Maarten: will check if the network team could have communicated the outage (OTG:0028285) better
        • CERN EOS issue (GGUS:119410): Upload files to CERN-BUFFER
        • Failed transfers: (GGUS:119448)
      • T1:

Sites / Services round table:

  • ASGC:
  • BNL: ntr
  • CNAF:
  • FNAL: ntr
  • GridPP:
  • IN2P3: ntr
  • JINR:
  • KISTI: ntr
  • KIT:
  • NDGF: ntr
  • NL-T1: ntr
  • NRC-KI:
  • OSG:
    • can GGUS:118799 be closed?
      • done
    • GGUS:117903 is an example of ~20 tickets about storage consistency checks for ATLAS:
      • opened in Nov, not making progress, what shall we do with them?
      • David C: will have a look from the ATLAS side
      • Michael: these are being followed up in US-ATLAS operations meetings and, when needed, ADC weekly meetings
  • PIC:
  • RAL: ntr
  • TRIUMF:

  • CERN batch and grid services: ntr
  • CERN storage services:
    • yesterday EOSALICE was down 15:20-20:00 CET
      • the name space crashed during a compactification, had to be repaired and restarted
  • Databases:
    • the CMSR DB firewall changes failed and had to be rolled back
      • the downtime was 20 min
      • the intervention will be rescheduled after resolution
    • the ALICE online DB was taken down on request from Mon 19:00 to Wed 16:00
      • there was a power cut on Tue
  • GGUS:
  • Grid Monitoring:
  • MW Officer:

AOB:

Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r8 - 2016-02-11 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback